# Zero-free regions for the independence polynomial on restricted graph classes

### v1

@@ Abstract.
Generalising the Heilman-Lieb Theorem from statistical physics, Chudnovsky and Seymour [J. Combin. Theory Ser. B, 97(3):350–357] showed that the univariate independence polynomial of any claw-free graph has all of its zeros on the negative real line. In this paper, we show that for any fixed subdivded claw $H$ and any $\Delta$, there is an open set $F\subseteq\mathbb{C}$ containing $0,\infty)$ such that the independence polynomial of any $H$-free graph of maximum degree $\Delta$ has all of its zeros outside of $F$. We also show that no such result can hold when $H$ is any graph other than a subdivided claw or if we drop the maximum degree condition.
@@ 1. Introduction
A simplicial clique $K$ in a graph $G$ is a clique such that for every vertex $v\in V(K)$, the neighbours of $v$ outside $K$ induce a clique in $G$, that is $G[\partial v\setminus V(K)]$ is a clique for all $v\in V(K)$. Write $\mathcal{G}^{\text{Cl-S}}$ for set of claw-free graphs for which every connected component has a simplicial clique and write $\mathcal{G}^{\textup{Cl-S}}_{k}$ for the set of graphs in $\mathcal{G}^{\textup{Cl-S}}$ whose largest clique has size at most $k$. The class $\mathcal{G}^{\textup{Cl-S}}$ has been previously studied, e.g., in [15, 12, 2]. It is also worth noting that given any claw-free graph of bounded degree say $\Delta$, one can delete at most $\Delta-1$ vertices (e.g. delete all but one neighbours of any vertex) to obtain a graph in $\mathcal{G}^{\textup{Cl-S}}$; thus every claw-free graph of bounded degree is very close to a graph in $\mathcal{G}^{\textup{Cl-S}}$. We prove the following in Section 4.
@@
For certain classes of trees, Bencs [4] showed real rootedness of the independence polynomial. The class of general trees of bounded degree has received considerable attention because, as it turns out, the zeros of trees of maximum degree $\Delta$ coincide exactly with zeros of general graphs of maximum degree $\Delta$ (see e.g. [4]). In this direction, Dobrushin [18, 19] and Shearer [32] independently established a zero-free disc for graphs of maximum degree $\Delta$ that is centered at zero of radius $\lambda_{s}(\Delta)=(\Delta-1)^{\Delta-1}/\Delta^{\Delta}$. While this zero-free disc cannot be enlarged in the negative direction, in the positive real direction, Peters and Regts [31] found a zero-free region that contains a neighbourhood of the interval $0,\lambda_{c}(\Delta))$, where $\lambda_{c}(\Delta)=(\Delta-1)^{\Delta-1}/(\Delta-2)^{\Delta}>\lambda_{s}(\Delta)$ thus resolving a conjecture of Sokal [33]. This began a sequence of works which together show that, for graphs of maximum degree $\Delta$, there is a cardioid region $C_{\Delta}\subseteq\mathbb{C}$ (that intersects $\mathbb{R}$ in the interval ($\lambda_{s}(\Delta)$, $\lambda_{c}(\Delta)$)) outside of which zeros of the independence polynomial are dense [9, 8] but inside of which there are both zeros [11] and zero-free regions [7] suggesting an exact description might be difficult to give (although there is an understanding of the limiting behaviour [6]).
@@ 2. Notation
For a graph $G=(V,E)$ and a vertex $v\in V$, we write $\partial v$ for the neighbours of $v$ in $G$. For $U\subseteq V$, we write $\partial U=\{v\in V\setminus U:v\in\partial u\text{ for some }u\in U\}=(\cup_{u\in U}\partial u)\setminus U$ for the boundary of $U$. We write $G[U]$ for the graph induced by $G$ on $U$. We write $U^{(\leq r)}:=\{S\subseteq U:|S|\leq r\}$.
@@ 3. Zero-free regions for graphs with no induced subdivided claw
The introduction of admissible pairs is motivated by a process used in [24] to sample red-blue pairs of random independent sets in $G$. The union $R\cup B$ of independent sets $R$ and $B$ induces connected components in the graph $G$. The inductive definition of admissible pairs given above mirrors the process of growing the red-blue connected component from a distinguished vertex $u$, presented in [24, Algorithm 1].
@@
We prove the lemma by induction on $|U|+|G|$. For the base case, when $(L,U)=(L,\emptyset)$, we have $Z(U)=Z(U\setminus\partial L)=1$ and (i) and (ii) immediately follow.
@@
We are given positive integers $t,\Delta$. Let $r=2\mathop{\mathrm{vol}}(\Delta,2t)$ and let $F=\{\lambda\in\mathbb{C}:\mathop{\mathrm{Re}}\nolimits(\lambda)\geq 0\text{ and }\lambda^{r}\in R(2^{\Delta r})\}$. For any $S_{t,t,t}$-free graph $G=(V,E)$ of maximum degree $\Delta$ and any $\lambda\in F$, setting $\lambda_{v}=\lambda$ for all $v\in V$, we have $\lambda_{S}\in R(2^{\Delta r})$ for any $S\in V^{(\leq r)}$. Thus we can apply Lemma 9 to conclude $Z_{G,\lambda}\not=0$. ∎
@@
$F\supseteq F^{*}=\left\{(x+\mathrm{i}y):|y|\leq 2^{-\frac{1}{2}(\Delta+3)r}x^{1-\frac{1}{2}r},x\geq 2^{-\Delta}\right\}.$
To see this, note that if $z=x+\mathrm{i}y\in F^{*}$ then $x\geq 2^{-\Delta}\geq 2|y|$ so
$|\mathop{\mathrm{Im}}\nolimits(z^{r})|=|\mathop{\mathrm{Im}}\nolimits((x+\mathrm{i}y)^{r})|=\left|\sum_{t\text{ odd}}^{r}\binom{n}{t}(\mathrm{i}y)^{t}x^{r-t}\right|\leq\sum_{t\text{ odd}}^{r}\binom{n}{t}|y|x^{r-1}\leq 2^{r}|y|x^{r-1},$
and $|\mathop{\mathrm{Re}}\nolimits(z^{r})|\geq(x-|y|)^{r}\geq(\frac{1}{2}x)^{r}$. Now it is easy to check that $z^{r}\in R(2^{\Delta r})$. Note that close to the origin, we know the disc of radius $(\Delta-1)^{\Delta-1}/\Delta^{\Delta}\approx e/\Delta$ is zero-free (see the Introduction) and this disc intersects $F^{*}$.
@@
Without loss of generality, assume $\lambda_{0}\geq 1$. Fix $\varepsilon>0$ and choose $\psi>0$ such that the bounded sector
$F=\{z:|z|<\lambda_{0}+\varepsilon\text{ and }|\arg z|<\psi\}$
is such that $\{z^{r}:z\in F\}\subseteq R(2^{\Delta r})$. Then, by Lemma 9, $F$ is a multivariate zero-free region for $Z_{G,\boldsymbol{\lambda}}$. ∎
@@ 4. Claw-free graphs
We begin this section with some structural properties of claw-free graphs. Recall that a simplicial clique in a claw-free graph $G$ is a clique $K$ in $G$ such that for every vertex $v\in V(K)$, $G[\partial v\setminus K]$ is a clique. Recall that $\mathcal{G}^{\textup{Cl-S}}$ is the set of claw-free graphs whose every component has a simplicial clique and $\mathcal{G}^{\textup{Cl-S}}_{k}$ is the set of graphs in $\mathcal{G}^{\textup{Cl-S}}$ whose largest clique has size at most $k$. We will require the following (perhaps surprising) structural lemma, which we believe is new since the authors of [15] investigate structural properties of the graph class $\mathcal{G}^{\textup{Cl-S}}$ but do not mention that it is hereditary.
@@
We do induction on $|U_{Px}|+|G+u_{K}|$. If $|U_{Px}|=0$ then $Z(U_{Px})=Z(U_{Px}\setminus\partial x)=1$ and both (i) and (ii) immediately follow.
@@ 5.2. Blow-ups of cycles (dense)
Here we generalise the previous example to graphs of higher degree (and larger cliques) to show that Theorem 6 gives a zero-free region that is not quite best possible, but close.
By $C_{2n}[K_{s}]$ we denote the lexicographic product of a cycle of length $2n$ and a clique of size $s\geq 2$. Thus, the vertex set of $C_{2n}[K_{s}]$ is $2n]\times[s]$ and each vertex subset $\{i,i+1\}\times[s]$ (with artithmetic modulo $2n$) induces a $2s$-clique $K_{2s}$. There are no further edges. It is easy to check that $C_{2n}[K_{s}]$ is claw free; indeed, $G[K_{s}]$ is claw free for any claw-free graph $G$ and integer $s\geq 2$.
Although $C_{2n}[K_{s}]$ is not a line graph (of a simple graph), it is the line graph of a multigraph and so is amenable to the method of Asano contractions (see Theorem 4 and appendix). As the cliques that comprise $C_{2n}[K_{s}]$ have size $2s$, this approach yields $R(s^{2})$ as a multivariate zero-free region. (Refer to Theorem 4.) Alternatively, since any maximum clique in $C_{2n}[K_{s}]$ has size $2s$ and is simplicial, our Theorem 6 yields $R(2s)$ as a zero-free region. Below we see that Theorem 6 does not quite give the optimal zero-free region but comes close.
@@ 5.4. Blow-ups of cycles (sparse)
Consider the collection $\mathcal{C}=\{C_{2n}[2K_{1}]:n\in\mathbb{N}\}$ of blown-up cycles defined above. There is no open multivariate zero-free region for $\mathcal{C}$ that contains the entire real axis.
@@
Suppose $\varepsilon>0$. Define a weighting $\boldsymbol{\lambda}^{\varepsilon}:[2n]\times[2]\to\mathbb{C}$ of the vertices of $C_{2n}[2K_{1}]$ by
@@
where $W\in\mathbb{R}$ and $\varepsilon>0$ will be chosen presently. As before, by grouping equivalent independent sets, we can write down an equivalent (in the sense of preserving the partition function) weighting $\boldsymbol{\mu}^{\varepsilon}$ of the simple cycle $C_{2n}$:
$\mu^{\varepsilon}_{i}=\begin{cases}z_{\varepsilon}:=(W+1)(w_{\varepsilon}+1)-1=(2W+1)+(W+1)\varepsilon\mathrm{i},&\text{if $i$ is even;}\\ \bar{z_{\varepsilon}}:=(W+1)(\bar{w_{\varepsilon}}+1)-1=(2W+1)-(W+1)\varepsilon\mathrm{i},&\text{if $i$ is odd.}\end{cases}$
We then have $Z_{C_{2n}[2K_{1}],\boldsymbol{\lambda}^{\varepsilon}}=Z_{C_{2n},\boldsymbol{\mu}^{\varepsilon}}$.
With a view to obtaining a contradiction, assume that $S$ is an open multivariate zero-free region containing the positive real axis. Choose $\varepsilon>0$ such that the interval $1+\varepsilon\mathrm{i},1-\varepsilon\mathrm{i}]$ is contained in $S$. Then choose $W$ so that $\varepsilon^{2}(W+1)>2$. With these choices, $z_{\varepsilon}$ and $\bar{z_{\varepsilon}}$ lie outside the parabola $R(1)$, i.e., outside the zero-free region for simple cycles. There are just a countable set of choices for the cycle $C_{2n}$, so we cannot expect $Z_{C_{2n},\boldsymbol{\mu}^{\varepsilon}}$ to be zero for any weighting exactly of the form $\boldsymbol{\mu}^{\varepsilon}:[2n]\to\{z_{\varepsilon},\bar{z}_{\varepsilon}\}$. However, we can find a nearby zero. As in the proof of Lemma 13, we can choose $\delta\in(0,\varepsilon)$ such that $\delta^{2}(W+1)>2$ continues to hold, and such that there exist $n\in\mathbb{N}$ and $\boldsymbol{\mu}^{\delta}:[2n]\to\{z_{\delta},\bar{z_{\delta}}\}$ such that $Z_{C_{2n},\boldsymbol{\mu}^{\delta}}=0$. As we noted earlier, this implies $Z_{C_{2n}[2K_{1}],\boldsymbol{\lambda}^{\delta}}=0$, contradicting the assumption that $S$ is zero free. ∎

### v2

@@ Abstract.
Generalising the Heilman-Lieb Theorem from statistical physics, Chudnovsky and Seymour [J. Combin. Theory Ser. B, 97(3):350–357] showed that the univariate independence polynomial of any claw-free graph has all of its zeros on the negative real line. In this paper, we show that for any fixed subdivided claw $H$ and any $\Delta$, there is an open set $F\subseteq\mathbb{C}$ containing $0,\infty)$ such that the independence polynomial of any $H$-free graph of maximum degree $\Delta$ has all of its zeros outside of $F$. We also show that no such result can hold when $H$ is any graph other than a subdivided claw or if we drop the maximum degree condition.
@@ 1. Introduction
A simplicial clique $K$ in a graph $G$ is a clique such that for every vertex $v\in V(K)$, the neighbours of $v$ outside $K$ induce a clique in $G$, that is $G[\partial v\setminus V(K)]$ is a clique for all $v\in V(K)$. Write $\mathcal{G}^{\text{Cl-S}}$ for set of claw-free graphs for which every connected component has a simplicial clique and write $\mathcal{G}^{\textup{Cl-S}}_{k}$ for the set of graphs in $\mathcal{G}^{\textup{Cl-S}}$ whose largest clique has size at most $k$. The class $\mathcal{G}^{\textup{Cl-S}}$ has been previously studied, e.g., in [13, 10, 2]. It is also worth noting (see Lemma 11) that given any connected claw-free graph of bounded degree say $\Delta$, one can delete at most $\Delta+1$ vertices [^2] Given any claw-free graph $G$, we claim that removing any vertex $v$ and all its neighbours gives a graph $G^{\prime}$ in $\mathcal{G}^{\text{Cl-S}}$. Indeed, each component $C$ of $G^{\prime}$ is adjacent to some neighbour $u$ of $v$ and it is easy to check that the neighbours of $u$ in $C$ form a simplicial clique (using that $G$ is claw-free). to obtain a graph in $\mathcal{G}^{\textup{Cl-S}}$; thus every claw-free graph of bounded degree is very close to a graph in $\mathcal{G}^{\textup{Cl-S}}$. We prove the following in Section 4.
@@
For certain classes of trees, Bencs [6] showed real rootedness of the independence polynomial. The class of general trees of bounded degree has received considerable attention because, as it turns out, the zeros of trees of maximum degree $\Delta$ coincide exactly with zeros of general graphs of maximum degree $\Delta$ (see e.g. [6]). In this direction, Dobrushin [18, 19] and Shearer [32] independently established a zero-free disc for graphs of maximum degree $\Delta$ that is centered at zero of radius $\lambda_{s}(\Delta)=(\Delta-1)^{\Delta-1}/\Delta^{\Delta}$. While this zero-free disc cannot be enlarged in the negative direction, in the positive real direction, Peters and Regts [31] found a zero-free region that contains a neighbourhood of the interval $0,\lambda_{c}(\Delta))$, where $\lambda_{c}(\Delta)=(\Delta-1)^{\Delta-1}/(\Delta-2)^{\Delta}>\lambda_{s}(\Delta)$ thus resolving a conjecture of Sokal [33]. This began a sequence of works which together show that, for graphs of maximum degree $\Delta$, there is a cardioid region $C_{\Delta}\subseteq\mathbb{C}$ (that intersects $\mathbb{R}$ in the interval ($\lambda_{s}(\Delta)$, $\lambda_{c}(\Delta)$)) outside of which zeros of the independence polynomial are dense [16, 8] but inside of which there are both zeros [9] and zero-free regions [5] suggesting an exact description might be difficult to give (although there is an understanding of the limiting behaviour as $\Delta\to\infty$ [4]).
@@ 2. Notation
For a graph $G=(V,E)$ and a vertex $v\in V$, we write $\partial v$ for the neighbours of $v$ in $G$. For $U\subseteq V$, we write $\partial U=\{v\in V\setminus U:v\in\partial u\text{ for some }u\in U\}=(\cup_{u\in U}\partial u)\setminus U$ for the boundary of $U$. We write $G[U]$ for the graph induced by $G$ on $U$. We write $U^{(\leq r)}:=\{S\subseteq U:1\leq|S|\leq r\}$.
@@ 3. Zero-free regions for graphs with no induced subdivided claw
The introduction of admissible pairs is motivated by a process used in [25] to sample red-blue pairs of random independent sets in $G$. The union $R\cup B$ of independent sets $R$ and $B$ induces connected components in the graph $G$. The inductive definition of admissible pairs given above mirrors the process of growing the red-blue connected component from a distinguished vertex $u$, presented in [25, Algorithm 1].
@@
We prove the lemma by induction on $|U|+|V|$. For the base case, when $(L,U)=(L,\emptyset)$, we have $Z(U)=Z(U\setminus\partial L)=1$ and (i) and (ii) immediately follow.
@@
We are given positive integers $t,\Delta$. Set $r=2\mathop{\mathrm{vol}}(\Delta,2t)$. For each $\ell\in\mathbb{N}$, let $F_{\ell}=\{\lambda\in\mathbb{C}:\lambda^{\ell}\in R(2^{\Delta r})\}$, which is an open set containing $0,\infty)$. Let
$F=\{\lambda\in\mathbb{C}:\lambda^{\ell}\in R(2^{\Delta r})\text{ \> for all \> }\ell=1,\ldots,r\}=\bigcap_{\ell=1}^{r}F_{\ell},$
which is also an open set containing $0,\infty)$. For any $S_{t,t,t}$-free graph $G=(V,E)$ of maximum degree $\Delta$ and any $\lambda\in F$, setting $\lambda_{v}=\lambda$ for all $v\in V$, we have $\lambda_{S}\in R(2^{\Delta r})$ for any $S\in V^{(\leq r)}$. Thus we can apply Lemma 9 to conclude $Z_{G,\lambda}\not=0$. ∎
@@
$F\supseteq F^{*}=\left\{(x+\mathrm{i}y):|y|\leq Ax^{-\frac{r-2}{2}},x\geq\Delta^{-1}\right\}.$
for a suitably small choice of constant $A$ depending only on $r$ and $\Delta$ (recall that $r$ itself only depends on $\Delta$ and $t$). To see this, we must show that for any $z=x+\mathrm{i}y\in F^{*}$, we have $z^{\ell}\in R(2^{\Delta r})$ for each $\ell=1,\ldots,r$. Note first that
$|\mathop{\mathrm{Im}}\nolimits(z^{\ell})|=|\mathop{\mathrm{Im}}\nolimits((x+\mathrm{i}y)^{\ell})|=\left|\sum_{t\text{ odd}}^{\ell}\binom{\ell}{t}(\mathrm{i}y)^{t}x^{\ell-t}\right|\leq\sum_{t\text{ odd}}^{\ell}\binom{\ell}{t}|y|x^{\ell-1}\leq 2^{\ell}|y|x^{\ell-1},$
and similarly $|\mathop{\mathrm{Re}}\nolimits(z^{\ell})|\geq x^{\ell}-2^{\ell}|y|^{2}x^{\ell-2}\geq\frac{1}{2}x^{\ell}$, where both inequalities assume that $A$ is sufficiently small (take e.g. $A\leq 2^{-\frac{1}{2}(r+1)}\Delta^{-\frac{1}{2}r}$). Now one can verify that $\mathop{\mathrm{Im}}\nolimits(z^{\ell})^{2}\leq 2^{-\Delta r}\mathop{\mathrm{Re}}\nolimits{z^{\ell}}$ for sufficiently small $A$ (take e.g. $A\leq 2^{-\frac{1}{2}(\Delta r+2r+1)}$) showing that $z^{\ell}\in R(2^{\Delta r})$.
While $F^{*}$ does not contain the origin, we know the disc of radius $(\Delta-1)^{\Delta-1}/\Delta^{\Delta}\approx e/\Delta$ is zero-free (see the Introduction) and this disc intersects $F^{*}$.
@@
Given any region $R\subseteq\mathbb{C}$ and positive integer $p$, we define $R^{\times p}:=\{\prod_{i=1}^{q}z_{i}:1\leq q\leq p\text{ and }z_{1},\ldots,z_{q}\in R\}$. Also, we write $S(\rho,\theta):=\{z:|z|<\rho\text{ and }|\arg z|<\theta\}$ for the bounded sector of radius $\rho$ and angle $\theta$ and we write $B(\rho):=\{z:|z|<\rho\}$ for the disc of radius $\rho$ centered at 0.
@@
First note that for $\rho_{1}\geq 1$, we have $S(\rho_{1},\theta)^{\times p}=S(\rho_{1}^{p},p\theta)$ and for $\rho_{2}\leq 1$, we have $B(\rho_{2})^{\times p}=B(\rho_{2})$. Furthermore, again assuming $\rho_{2}\leq 1\leq\rho_{1}$, it is easily checked that $(S(\rho_{1},\theta)\cup B(\rho_{2}))^{\times p}\subseteq S(\rho_{1}^{p},p\theta)\cup B(\rho_{1}^{p-1}\rho_{2})$.
In the statement of Theorem 5, we are given $\lambda_{0}>0$ and positive integers $t,\Delta$. As before, set $r=2\mathop{\mathrm{vol}}(\Delta,2t)$ and without loss of generality, we may assume $\lambda_{0}\geq 1$. Fix $\varepsilon>0$ and set $\rho_{1}=\lambda_{0}+\varepsilon$. From the discussion in the previous paragraph, we can choose $\theta,\rho_{2}>0$ sufficiently small such that the set
$F:=S(\rho_{1},\theta)\cup B(\rho_{2})$
satisfies that $F^{\times r}\subseteq R(2^{\Delta r})$. Then, by Lemma 9, $F$ is a multivariate zero-free region for $Z_{G,\boldsymbol{\lambda}}$ for any $S_{t,t,t}$-free graph of maximum degree $\Delta$. ∎
@@ 4. Claw-free graphs
We begin this section with some structural properties of claw-free graphs. Recall that a simplicial clique in a claw-free graph $G$ is a clique $K$ in $G$ such that for every vertex $v\in V(K)$, $G[\partial v\setminus K]$ is a clique. Recall that $\mathcal{G}^{\textup{Cl-S}}$ is the set of claw-free graphs whose every component has a simplicial clique and $\mathcal{G}^{\textup{Cl-S}}_{k}$ is the set of graphs in $\mathcal{G}^{\textup{Cl-S}}$ whose largest clique has size at most $k$. We will require the following structural lemma; similar results are mentioned in [15, 10].
@@
We do induction on $|U_{Px}|+|V|$. If $|U_{Px}|=0$ then $Z(U_{Px})=Z(U_{Px}\setminus\partial x)=1$ and both (i) and (ii) immediately follow.
@@ 5.2. Blow-ups of cycles (dense)
Here we generalise the previous example to graphs of higher degree (and larger cliques) to show that Theorem 6 gives a zero-free region that is close to best possible.
By $C_{2n}[K_{s}]$ we denote the lexicographic product of a cycle of length $2n$ and a clique of size $s\geq 2$. Thus, the vertex set of $C_{2n}[K_{s}]$ is $2n]\times[s]$ and each vertex subset $\{i,i+1\}\times[s]$ (with arithmetic modulo $2n$) induces a $2s$-clique $K_{2s}$. There are no further edges. It is easy to check that $C_{2n}[K_{s}]$ is claw free; indeed, $G[K_{s}]$ is claw free for any claw-free graph $G$ and integer $s\geq 2$.
Although $C_{2n}[K_{s}]$ is not a line graph (of a simple graph), it is the line graph of a multigraph and so is amenable to the method of Asano contractions (see Theorem 4 and appendix). As the cliques that comprise $C_{2n}[K_{s}]$ have size $2s$, this approach yields $R(s^{2})$ as a multivariate zero-free region. (Refer to Theorem 4.) Alternatively, since any maximum clique in $C_{2n}[K_{s}]$ has size $2s$ and is simplicial, our Theorem 6 yields $R(2s)$ as a zero-free region. The result below shows that Theorem 6 is within a factor of $2$ of being best possible (i.e. we cannot replace $R(k)$ with $R(k^{\prime})$ in the statement if $k^{\prime}<k/2$).
@@ 5.4. Blow-ups of cycles (sparse)
Consider the collection $\mathcal{C}=\{C_{2n}[2K_{1}]:n\in\mathbb{N}\}$ of blown-up cycles defined above. There is no open multivariate zero-free region for $\mathcal{C}$ that contains the positive real axis.
@@
Given constants $\varepsilon>0$ and $W>0$, define a weighting $\boldsymbol{\lambda}^{\varepsilon,W}:[2n]\times[2]\to\mathbb{C}$ of the vertices of $C_{2n}[2K_{1}]$ by
@@
As before, by grouping equivalent independent sets, we can write down an equivalent (in the sense of preserving the partition function) weighting $\boldsymbol{\mu}^{\varepsilon,W}$ of the simple cycle $C_{2n}$:
$\mu^{\varepsilon,W}_{i}=\begin{cases}z_{\varepsilon,W}:=(W+1)(w_{\varepsilon}+1)-1=(2W+1)+(W+1)\varepsilon\mathrm{i},&\text{if $i$ is even;}\\ \bar{z}_{\varepsilon,W}:=(W+1)(\bar{w}_{\varepsilon}+1)-1=(2W+1)-(W+1)\varepsilon\mathrm{i},&\text{if $i$ is odd.}\end{cases}$
We then have $Z_{C_{2n}[2K_{1}],\boldsymbol{\lambda}^{\varepsilon,W}}=Z_{C_{2n},\boldsymbol{\mu}^{\varepsilon,W}}$.
With a view to obtaining a contradiction, assume that $S$ is an open multivariate zero-free region for $\mathcal{C}$ that contains the positive real axis. Choose $\delta>0$ such that the interval $1-\delta\mathrm{i},1+\delta\mathrm{i}]$ is contained in $S$. We will check that there exists $W>0$, $\varepsilon\in(0,\delta]$, and $n\in\mathbb{N}$ such that $Z_{C_{2n},\boldsymbol{\mu}^{\varepsilon,W}}=0$ and hence $Z_{C_{2n}[2K_{1}],\boldsymbol{\lambda}^{\varepsilon,W}}$ giving the contradiction.
Consider the set
$T_{\delta}=\{z_{\varepsilon,W}:\varepsilon\in(0,\delta],\;W>0\}$ $=\{(2W+1)+(W+1)\varepsilon\mathrm{i}:\;\varepsilon\in(0,\delta],\;W>0\}$
$=\{x+\mathrm{i}y:0<y<\tfrac{1}{2}\delta(x+1),\;x>1\}.$
It is easy to check that $T_{\delta}$ contains a (unbounded) part of the parabola $\partial R(1)$. Lemma 13 then implies that there exists $z_{\varepsilon,W}\in T_{\delta}$ and $n\in\mathbb{N}$ such that alternately weighting the vertices of $C_{2n}$ with $z_{\varepsilon,W}$ and $\bar{z}_{\varepsilon,W}$ gives a zero of $Z_{C_{2n}}$, i.e. $Z_{C_{2n},\boldsymbol{\mu}^{\varepsilon,W}}=0$. ∎